OpenStudy (anonymous):

A particular circle in the standard (x,y) coordinate plane has an equation of (x-5) squared + y squared = 38. What are the radius of the circle, in coordinate units, and the coordinates of the center of the circle?

OpenStudy (anonymous):

The coordinates of the centre are (-1,0)

OpenStudy (anonymous):

Could you give me an explanation of how you got your answer?

OpenStudy (anonymous):

Also the + was actually supposed to be an = so excuse that my bad

OpenStudy (anonymous):

expand the equation given to get \[x ^{2}+y ^{2}-2x +63 = 0\]

OpenStudy (anonymous):

general equation of a circle is x^2 + Y^2 +2gx + 2fy + c = 0

OpenStudy (anonymous):

The equation is \[(x-5)^{2}+y ^{2}=38\]

OpenStudy (anonymous):

Comparing the two we get g = -1 y=0 Also centre is equal to (-g.-f)

OpenStudy (anonymous):

sorry the centre would be (1,0)

OpenStudy (anonymous):

I think the center would be 5,0 because if we set x to =0 we get x=5 and since the y is only squared the y would automatically be 0. I just can't remember the radius. All i know is that it's the number the whole thing is equal to and i think it's square rooted.

OpenStudy (anonymous):

Thanks though! I think i've got the answer :)